OpenStudy (anonymous):

Jennie has a jar of 54 coins consisting of dimes and quarters. If she has a total of $11.10, how many dimes does she have?

OpenStudy (anonymous):

d + q = 54 d(10) + q(25) = 1110 Do you know substitution?

OpenStudy (anonymous):

no not really

OpenStudy (anonymous):

Okay, here are the steps then. First find q(quarters) in the first equation. d + q = 54 Subtract d. q = 54 - d Now put the value of q into the second equation. 10d + 25q = 1110 10d + 25(54 - d) = 1110 Distribute. Solve for d(dimes) 10d + 25(54) + 25(-d) =1110 10d + 79 - 25 = 1110 10d + 54 = 1110 Can you solve from there? 10d + 54 = 1110

OpenStudy (anonymous):

Is it 16 or 22?

OpenStudy (anonymous):

so it is 16?

OpenStudy (anonymous):

One sec...

OpenStudy (anonymous):

10d + 25(54) + 25(-d) =1110 10d + 1350 + 25(-d) =1110 @jazy thats the mistake i caught

OpenStudy (anonymous):

Okay! Thanks @sudawi Yes you're right @Eulie it is 16. 10d + 25(54) + 25(-d) =1110 10d + 1350 + 25(-d) =1110 10d + 1350 - 25d = 1110 -15d + 1350 = 1110 Subtract 1350. -15d = -240 Divided by -15 d = 16